OpenStudy (anonymous):

What is 30% from $39.50

OpenStudy (anonymous):

$11.85

OpenStudy (amistre64):

39.50 - 39.50(.3)

OpenStudy (amistre64):

27.65 would be the 'sale' price

OpenStudy (anonymous):

how did you get that answer amistre and cocople

OpenStudy (anonymous):

because both of your answers is a choie on my paper

OpenStudy (anonymous):

i meant choice

OpenStudy (amistre64):

........... I got the answer by taking the price given (39.50) and subtracting 30% of it away -(39.50*.3) but its all in how you interpret the question

OpenStudy (anonymous):

30/100 x 39.50

OpenStudy (anonymous):

39.50 X 30/100 = 11.85

OpenStudy (anonymous):

ok how did you sub. it?

OpenStudy (amistre64):

what does 'sub. it' mean?

OpenStudy (anonymous):

subtract

OpenStudy (amistre64):

i subed it the same way I posted to begin with 39.50 - 39.50(.3) = 'sale' price

OpenStudy (amistre64):

39.50 - 39.50(.3) = 'sale' price <- equals the price you pay ^^ ^ ^^ ^^ ^ 30% of original price ^^ minus original price

OpenStudy (amistre64):

percents are just fractions; you have part of something --------------- = decimal value of percentage all of it

OpenStudy (amistre64):

part of it ------- = .30 39.50 part of it = .30(39.50)

OpenStudy (anonymous):

Yes work out 1% by dividing 39.50 by 100 then multiply by the percentage you need in this case 30%. That's the easiest way for me to remember it otherwise its confusing as to which number to be at the bottom etc. So its\[\frac{39.50}{100} \times 30\] now 39.50−11.85=27.65

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

also allows you to do it in your head. if you move the decimal two places left you get 1% then multiply by the percentage. The division is more difficult to do in your head.

OpenStudy (amistre64):

yeah, thats why i use an abacus ;)

OpenStudy (anonymous):

ha ha ha ha, nice

OpenStudy (anonymous):

WAit i got to different answer wich one is correct 27.65 11.85 32.50 39.20

OpenStudy (anonymous):

11.85

OpenStudy (amistre64):

pick one and if its right; you know which answer they want; and if its wrong, you still know which answer they are looking for :)

OpenStudy (anonymous):

27.65 because it says 30% FROM 39.50 = 27.65 11.85 is the 30% the other two are just wrong!

OpenStudy (anonymous):

just ask Google

OpenStudy (anonymous):

wow XD